OpenStudy (anonymous):

Please help! Show me how to factor these binomials! m^4-256 and z^3+125, thank you! Descriptions are appreciated :)

OpenStudy (anonymous):

m^4-256 this is a perfect square

OpenStudy (anonymous):

because 2*2=4, right?

OpenStudy (anonymous):

the 1 you want to concentrate on is 256

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

try and get it into the form \(a^2-b^2\) Use the hints that mth3v4 has given you.

OpenStudy (anonymous):

cool pic asna :D

OpenStudy (asnaseer):

thx :) I'm actually an alien from another planet :D

OpenStudy (anonymous):

:)

OpenStudy (asnaseer):

Starshine7 - can you express 256 as (something)^2 ?

OpenStudy (anonymous):

16x16?

OpenStudy (anonymous):

:)

OpenStudy (asnaseer):

good, now can you express \(m^4\) as (something)^2 ?

OpenStudy (anonymous):

so can you put that in in a factored form

OpenStudy (anonymous):

m*m*m*m?

OpenStudy (asnaseer):

Yes, so \(m^4=(m^2)^2\) Now use those two to get the expression into the form \(a^2-b^2\) so that you can then factor it like this:\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

why is it written like (m+number)(trinomial)? final answer i mean. If it helps, I factor by grouping if i can incorporate that somehow

OpenStudy (asnaseer):

just do this one step at a time and al will be revealed :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lol oops

OpenStudy (anonymous):

nothing happened

OpenStudy (asnaseer):

:)

OpenStudy (asnaseer):

Starshine7 - are you stuck?

OpenStudy (anonymous):

Sadly yes :(

OpenStudy (anonymous):

look at the first answer you gave me

OpenStudy (anonymous):

how did you come up with that ?

OpenStudy (asnaseer):

ok, we saw above that:\[256=16^2\]\[m^4=(m^2)^2\]so now we can rewrite your original equation as:\[m^4-256=(m^2)^2-16^2\]now mak use of:\[a^2-b^2=(a+b)(a-b)\]to simplify this further.

OpenStudy (anonymous):

after that if can be factored again as an option as you can see

OpenStudy (anonymous):

(m^2)^2 confuses me

OpenStudy (anonymous):

(i think i am confusing starshine)

OpenStudy (asnaseer):

ok, lets use a substitution to make it clearer...

OpenStudy (asnaseer):

let \(n=m^2\), then we have:\[m^4-256=(m^2)^2-16^2=n^2-16^2\]

OpenStudy (anonymous):

why n? o.O

OpenStudy (asnaseer):

we could use any other symbol - it doesn't really matter

OpenStudy (asnaseer):

the point is it is now in the form \(a^2-b^2\)

OpenStudy (asnaseer):

so you should now be able to factorise it.

OpenStudy (anonymous):

okies properties of exponents state when you multiply exponent and exponent you add the exponent values together (a^x)(a^b)=a^x+b

OpenStudy (anonymous):

exponent properties/laws of exponents whatever

OpenStudy (asnaseer):

Starshine7 - are you familiar with the exponent laws mth3v4 has shown? i.e.:\[(x^a)^b=x^{ab}\]

OpenStudy (anonymous):

If it helps I dont understand when a and b are examples, I was taught using only the numbers, so thats why im really confused :(

OpenStudy (anonymous):

those are just any number

OpenStudy (anonymous):

I know >_> Im weird like that

OpenStudy (asnaseer):

so you are comfortable with something like this:\[18^2-16^2=(18+16)(18-16)\]but when replaced with symbols like this:\[a^2-b^2=(a+b)(a-b)\]you get lost?

OpenStudy (anonymous):

Yes, exactly :)

OpenStudy (asnaseer):

ok - the 'a' and 'b' are just placeholders for any number

OpenStudy (anonymous):

I just have attention issues so the numbers help me understand it better, i drift easily.

OpenStudy (asnaseer):

what this says is that the formula:\[a^2-b^2=(a+b)(a-b)\]is true for ANY value of 'a' and 'b'

OpenStudy (asnaseer):

so in your example above, we ended up with:\[m^4-256=(m^2)^2-16^2=n^2-16^2\]which means we can think of it as a=n and b=16 to get:\[n^2-16^2=(n+16)(n-16)\]we can then substitute back \(n=m^2\) to get:\[m^4-256=(n+16)(n-16)=(m^2+16)(m^2-16)\]

OpenStudy (asnaseer):

do you understand those steps?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

(a ^x ) ( a ^b )= a^x+b ^ ^ ^ ^ ^ ^ ^ | | | | / | | | L _____|___ L _ |___L_ _ L _ exponent number both x and b can be any number | | | your number ______ |

OpenStudy (anonymous):

does my funny drawing help XD

OpenStudy (asnaseer):

ok, so now you should notice that we have \((m^2-16)\) as one of the factors. since \(16=4^2\), we can write this as:\[(m^2-16)=(m^2-4^2)=(m+4)(m-4)\]using the same formula. we therefore end up with:\[m^4-256=(m^2+16)(m+4)(m-4)\]

OpenStudy (anonymous):

So you just keep factoring to the lowest numbers that are perfect squares?

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

mth3v4 - you have a lot of patience to draw all that :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

it still comes out a bit messy

OpenStudy (anonymous):

(a^x) (a^y) = a ^ x+y lemme try this you have this thing here right starshine?

OpenStudy (asnaseer):

Starshine7 - for your second expression, you need to use the following factorisation:\[a^3+b^3=(a+b)(a^2-ab+b^2)\]so you need to get your expression:\[z^3+125\]into the form \(a^3+b^3\)

OpenStudy (anonymous):

okies finish what you have first

OpenStudy (anonymous):

I guess

OpenStudy (asnaseer):

I have to go now guys - Starshine I am sure mth3v4 will help you with the rest. Good luck all...

OpenStudy (anonymous):

Aww..well bye Asna :) thanks anyways

OpenStudy (anonymous):

kk bb

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

z^3+125 is your next 1

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

it is not going to be squared but cubed

OpenStudy (anonymous):

have you heard of diff of cubes or sum of cubes formulas

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so what ansa wrote there before leaving was formula for diff of cubes

OpenStudy (anonymous):

because you just see just a 2 termed expression here

OpenStudy (anonymous):

so how would you make x^3 and change it to x^2

OpenStudy (anonymous):

better yet lemme ask the question another way how do you get x^3 from x^2

OpenStudy (anonymous):

I'm not sure. I think they only have 2 exs in common

OpenStudy (anonymous):

okies lets try this ( a^x ) ( a^y ) = a ^ x +y the variable "a" can be any number 1, 2, 3, whatever, happy face , i dont care what you want it to be ok so far

OpenStudy (anonymous):

this just goes fot the variable "a" we good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then next this for the variables "x, y" these give a numeric value (these have to be a number) giving a value ( a^x ) ( a^y) = a ^ x + y

OpenStudy (anonymous):

so in the end the numeric value of your "a" exponent value "x" "a" exponent value "y" they add up result for the a ^ x + y is this more understandable?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so can we try that again how to get x^3 from x^2

OpenStudy (anonymous):

(x)(x+x)?

OpenStudy (anonymous):

i just want it from x^2

OpenStudy (anonymous):

I'm sorry I have no idea what to do :(

OpenStudy (anonymous):

since it is already squared right (x^2)(x)=x how many more "x" do we need for x ^3

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

you have x^2 already

OpenStudy (anonymous):

but x^3 is made up of x^3 theres not anything left to do

OpenStudy (anonymous):

remember (x^2)(x)=x? ^ | value of exponent you are adding to the result

OpenStudy (anonymous):

lol if there is nothing then say nothing

OpenStudy (anonymous):

be more sure of you self now just give the reason

OpenStudy (anonymous):

nm that its probably more difficult to explain

OpenStudy (anonymous):

but do you understand how it works

OpenStudy (anonymous):

thats all i want to know

OpenStudy (anonymous):

How about we start over with a new problem, m^3+216n^3 Show me in steps how to solve this and ill try another on my own, this might work

OpenStudy (anonymous):

okeedokee